OpenStudy (anonymous):

Square root of 3 times square root of 21????

OpenStudy (agreene):

\[\sqrt{a}\sqrt{b}=\sqrt{a*b}\] so, \[\sqrt{3*21}\] evaluate that

OpenStudy (anonymous):

Ohh ok thanks!!! @agreene

OpenStudy (agreene):

no problem